A computer monitor has a height of 9 inches with a diagonal measurement of 15 inches. How many inches wide is the rectangular screen?
Answer by Jolliano(16)

In a rectangle,the height,width and diagonal form a right angled triangle with the diagonal as the hypotenuse.
Using Pythagoras theorem,
a^2 + b^2 = c^2
w^2 + 9^2 = 15^2
w^2 + 81 = 225
w^2 = 225 - 81
w^2 = 144
w = 12 inches